# Supplement to “Steering LLMs Towards Unbiased Responses: A Causality-Guided Debiasing Framework”

### v1

@@ 4.2.2 Our Results (Discrim-Eval)
We further examine the relative gaps for specific demographic features. For example, in Figure 6, focusing on GPT-3 (which has the smallest relative gaps among the 4 chosen LLMs), we observe a growing trend of relative gaps on the default response (black bars) as the age increases. This suggests that GPT-3’s default responses may be more biased towards more senior individuals compared to junior ones. It is worth noting that the trend may potentially differ across LLMs, and we include additional plots for other LLMs on different demographic groups in Appendix D.2.
5 Conclusion
This paper presents a causality-guided and prompting-based LLM debiasing framework. In particular, we highlight the critical role of selection mechanisms in modeling data corpus bias and, more importantly, formulating how prompt designs can influence LLM outputs by specifying different selection conditions on its internal representations. Guided by causal understandings of such interplay, we identify principled prompting strategies for more effective debiasing. Our strong empirical results demonstrate the benefits of our framework, offering clear intuitions and theoretical foundations for effective debiasing approaches. Future work will naturally extend to acquiring bias-free knowledge and representations for LLMs.
Broader Impact
In this paper, we present a causality-based LLM output debiasing framework. We aim to provide causal understandings on both the underlying generating process for the training data corpus, and the LLM reasoning process where the output is modulated by input prompts through selection mechanisms. Our proposed framework serves as a general solution to promote fairness in pretrained LLMs without accesssing their model parameters. This framework is applicable to many socially important domains, and we do not foresee any particular potential negative social impact or ethical concerns.
Acknowledgement
This material is based upon work supported by the AI Research Institutes Program funded by the National Science Foundation (NSF) under AI Institute for Societal Decision Making (AI-SDM), Award No. 2229881. This project is also partially supported by an Amazon Research Award (Fall 2022 CFP), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Apple Inc., KDDI Research Inc., Quris AI, and Infinite Brain Technology.
Appendix A Related Works
In this section, we provide detailed discussions on related works. We consider the combinations of very related topics including causality, algorithmic fairness, and LLM reasoning. In Section A.1, we consider causal notions of fairness that do not specifically pertain to the LLM context. In Section A.2, we consider existing efforts to draw connections between causality and LLM reasoning. In Section A.3, we consider previous works on LLM debiasing. In Section A.4, we consider previous works that involve all three topics.
A.1 Causality and Fairness
It has been recognized in the algorithmic fairness literature that causality provides a unique tool to facilitate a better understanding of the data-generating process, and therefore, more effective bias quantifications and mitigations (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019; Nabi et al., 2022; von Kügelgen et al., 2022; Tang et al., 2023a). Previous causal fairness literature has considered notions based on estimating or bounding various kinds of causal effects (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019), and also the causal modeling of the dynamics as well as the long-term implications of bias mitigation strategies (Creager et al., 2020; Zhang et al., 2020; Tang et al., 2023a, b).
A.2 Causality and LLMs
A.3 Debiasing Language Models
There is a large amount of work discussing bias and fairness in the context of language models (LMs) (Bordia & Bowman, 2019; Liang et al., 2021; Abid et al., 2021; Wang et al., 2023a; Liu et al., 2023; Ray, 2023; Rozado, 2023), and our investigation lies on debiasing techniques with causal understandings of the sources of biases. For debiasing approaches in the context of LMs, there are proposals involving direct fine-tuning of model parameters (Kaneko & Bollegala, 2021; Garimella et al., 2021; Lauscher et al., 2021; Guo et al., 2022), modifying the decoding steps (Schick et al., 2021), incorporating Reinforcement Learning with Human Feedback (RLHF) to better align the models with human values (Ouyang et al., 2022; Bai et al., 2022; Yao et al., 2023), and prompting-based techniques (Si et al., 2022; Tamkin et al., 2023; Oba et al., 2023; Ganguli et al., 2023). We focus on prompting-based techniques and identify principles for prompt designs to steer LLMs toward unbiased responses by a). reducing biased reasoning and b). encouraging bias-free reasoning. We provide demonstrations of how we can employ the above two principles. Works on LLMs reasoning such as Wei et al. (2022); Zheng et al. (2023) can also be incorporated into with our framework to encourage bias-free reasoning.
A.4 Debiasing LLMs from Causal Perspectives
Most closely related literature considers LLM debiasing strategies from causal perspectives. Vig et al. (2020) utilize causal mediation analysis and consider neuron-level intervention to investigate the instantiation of gender bias in Transformer-based language models (Vaswani et al., 2017). Zhou et al. (2023) propose Causal-Debias to mitigate the unwanted stereotypical association by fine-tuning pretrained language models. The causal model they consider involves four variables: label-relevant factor, bias-relevant factor, raw sentence, and ground-truth label. Wang et al. (2023b) pay special attention to entity bias and propose a specific structural causal model (SCM) for easier parameter estimations, such that the intervention-based mitigation strategy can be carried out. Their causal model involves four variables: entity, raw text, LLM input, and LLM output.
(a) LLM reasoning process (annotated with selection mechanisms specified by prompts)
(b) LLM reasoning process (Strategy I: nudge towards demographic-agnositic fact)
(c) LLM reasoning process (Strategy II: counteract existing selection bias)
(d) LLM reasoning process (Strategy III: nudge aware from demographic-aware text)
(e) LLM reasoning process (all strategies combined)
Figure 7: Additional illustrations on debiasing strategies.
Appendix B Further Illustrations and Discussions of Our Framework
In Figure 7, we use light coral (blue) to highlight unregulated (regulated) information flows from demographic representations to LLM outputs in the internal reasoning process. We use mix-colored highlight to denote partially regulated information flow, e.g., the edge from demographic-aware text representation to LLM potential output in Figure 6(d). Comparing Figure 6(e) (all strategies combined) with Figures 6(b)6(d) (strategies applied individually), we can see that collectively, the strategies address the social bias in language more comprehensively.
Here by “regulated”, we are referring to the constraint over the association between demographic information representation and the potential output. We would like to note that our causal models can naturally handle situations where certain kinds of neutral dependence patterns involving demographic information are not necessarily considered problematic (Tang et al., 2024). We present relevant experimental results in Section D.1. We would also like to note that our causality-guided framework is not limited to situations where demographic information is explicitly involved. One can adapt our framework to the specific practical setting by incorporating relevant nodes into causal graphs, thereby identifying the most suitable debiasing strategies therein.
Appendix C Experimental Details
Pretrained LLMs
For the GPT models used in our experiments, we consider snapshots from June 13th, 2023 where the knowledge cut-off time is Sep 2021. Since the legacy GPT-3 model (a.k.a., text-davinci-003) is no longer supported when we conduct the experiments, we use the model “gpt-3.5-turbo-instruct” instead as it has similar capabilities as GPT-3 era models. The Mistral-7B model we use in our experiments is the improved instruction fine-tuned version (a.k.a., “Mistral-7B-Instruct-v0.2”). We are unable to use Claude 2 for our experiments on the Discrim-Eval dataset because we cannot access the probabilities of generated tokens. All LLMs responses are obtained with a temperature of 0.
Iterative Prompting
C.1 Gender Bias: WinoBias
Since there are 571 examples where two or more models make the same mistakes on the factual questions. After looking into these mistakes, we find that some of the examples are indeed ambiguous in nature (e.g., ”The construction worker stared at the teacher because [who] was angry”). Therefore, we collected human evaluations to understand the inherent ambiguity in the WinoBias dataset. For each of the 571 examples, we collect 5 human responses on the factual questions (i.e., which sentence is more likely to happen in real life), and we ask the human annotator to choose from {Sentence A, Sentence B, and Equally likely}. Based on the human responses, we identified 60 examples (55 from the Type I task and 5 from the Type II task) where three or more annotators disagree with the ground truth answers or think both sentences in the factual questions are equally likely to happen in real life. We removed these 60 examples during our evaluation, and we will include the human evaluation results in our codebase.
C.2 Demographic Bias: Discrim-Eval
Table 3: Error Analysis on WinoBias Type II coreference task. We divide the models’ responses into 4 categories to better understand their success and failure cases. $T$ denotes the (%) of examples where LLM answers both the factual question and the original question correctly, $F$ denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge. TF denotes the coreference errors caused by gender bias as only factual questions are correctly answered, and FT indicates coreference success that may be due to gender shortcut since the factual questions get wrong but original questions are correctly answered.
GPT-4   Reduce + Fact 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25 Fact Only 99.36 99.87 0.25 0.00 0.13 0.00 0.25 0.13 Reduce Only 98.98 99.49 0.76 0.13 0.00 0.38 0.25 0.00 Default 98.73 99.75 0.89 0.13 0.25 0.00 0.13 0.13
Table 4: Error analysis of Reduce + Fact on Type I and Type II questions in WinoBias.
82.47 82.34 1.27 1.40 0.76 1.91 15.50 14.36   GPT-3.5   Type I 70.15 78.70 10.58 2.04 2.58 5.97 16.55 13.16 Type II 88.69 89.71 1.14 0.13 3.43 4.70 6.73 5.46   Claude   Type I 73.00 73.00 2.99 2.99 1.09 2.17 22.93 21.85 Type II 79.80 81.19 6.61 5.21 2.54 4.07 11.05 9.53   GPT-4   Type I 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Type II 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25
Table 5: Adjusting the levels of counteracting existing selection bias on GPT-3.5 with Reduce + Fact. We investigated how the level of counteracting stereotype impacts the model performance by adjusting the Fact part in the input prompt. For example, a counteract level of 100% indicates the following prompt: Assume that the physician can be male $0\%$ of the time and female $100\%$ of the time, and assume that the secretary can be male $0\%$ of the time and female $100\%$ of the time.
Appendix D Additional Experimental Results
D.1 Gender Bias: WinoBias
We include additional ablation studies on the WinoBias dataset, which include a detailed error analysis on WinoBias Type II coreference task (Table C.2), a detailed error analysis of Reduce + Fact with 4 LLMs on Type I and Type II tasks (Table C.2), and an ablation study on adjusting the levels of biased reasoning (Table C.2) in the prompt design.
D.2 Demographic Bias: Discrim-Eval
@@
In Figure 8, Figure 9, Figure 10, Figure 11, we show in details the performance comparison on Discrim-Eval across three demographic categories across four LLMs (Mistral 7B, GPT 3, GPT 3.5, and GPT 4). The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). The baseline setting is colored black.
Figure 8: Performance comparison on Discrim-Eval across three demographic categories in details on Mistral (7B). The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
@@
Figure 9: Performance comparison on Discrim-Eval across three demographic categories in details on GPT-3. The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
@@
Figure 10: Performance comparison on Discrim-Eval across three demographic categories in details on GPT-3.5. The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
@@
Figure 11: Performance comparison on Discrim-Eval across three demographic categories in details on GPT-4. The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.

### v2

@@ D.2.2 Our Results (Discrim-Eval)
We further examine the relative gaps for specific demographic features. For example, in Figure 6, focusing on GPT-3 (which has the smallest relative gaps among the 4 chosen LLMs), we observe a growing trend of relative gaps on the default response (black bars) as the age increases. This suggests that GPT-3’s default responses may be more biased towards more senior individuals compared to junior ones. It is worth noting that the trend may potentially differ across LLMs, and see additional plots for other LLMs on different demographic groups in Appendix D.2.3.
D.2.3 Additional Plots
In Figures 710, we show in detail the performance comparison on Discrim-Eval across three demographic categories across four LLMs (Mistral 7B, GPT-3, GPT-3.5, and GPT-4). The baseline setting is colored black. Across all four LLMs, we see that encouraging fact-based reasoning universally reduces the relative gap when used together with methods that discourage biased reasoning.
@@ D.2.3 Additional Plots
Figure 7: Performance comparison on Discrim-Eval across three demographic categories with Mistral (7B). The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning).
@@
Figure 8: Performance comparison on Discrim-Eval across three demographic categories with GPT-3. The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning).
@@
Figure 9: Performance comparison on Discrim-Eval across three demographic categories with GPT-3.5. The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning).
@@
Figure 10: Performance comparison on Discrim-Eval across three demographic categories with GPT-4. The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning).
[…]
[supp]